Ricky is 26 yrs old and Jenny is 5 yrs old. in how many yrs will ricky be twice as old as jenny?
Answer by ankor@dixie-net.com(22740)

: Ricky is 26 yrs old and Jenny is 5 yrs old.
in how many yrs will ricky be twice as old as jenny?
:
Let t = no. of yrs for this to be true
:
26 + t = 2(5+t)
26 + t = 10 + 2t
26 - 10 = 2t - t
16 = t
;
:
Check:
26 + 16 = 2(5 + 16)
42 = 2(21)